OpenStudy (sette):

Solve for k . i keep getting 11.1333333333 -20-6k=16-9(-k+19) k= ?

OpenStudy (anonymous):

\[-20-6k=16-9(-k+19)\]\[-20-6k=16+9k-171\]\[-20-16+171=9k-3k\]\[135=3k\]\[45=k\]